OpenStudy (marigirl):

The number of bacteria present in a culture at any time , t hours, can be modelled by the equation N=Noe^kt If the original number is doubled in 3 hours, find k

OpenStudy (danjs):

remember if the time t=0, the value will be No \[\large N(0)=N _{0}*e^0=N _{0}\]

OpenStudy (danjs):

They tell you the number has doubled at 3 hours, You will have 2*No, when t=3 \[\large N(3)=2*N _{0}=N _{0}*e^{3t}\]

OpenStudy (danjs):

sorry it is the k left in the exponent.. put in t=3 \[\large 2*N _{0}=N _{0}*e^{3k}\]

OpenStudy (danjs):

Solve for the k value, you get all that?

OpenStudy (marigirl):

a little lost @DanJS

OpenStudy (danjs):

ok. The exponential form they give you , \[\large N=N _{0}*e^{kt}\] Shows the number bacteria N, for any time t. The No is the original number bacteria at the start, when t=0. Putting in t=0 will show you, \[N=N _{0}*e^0 = N _{0}*1 = N _{0}\] you get that far?

OpenStudy (wolf1728):

N=No*e^kt Where No is beginning amount and N is ending amount where "e" is the mathematical constant 2.718281828 If No is (let's say 100 for example) Then after 3 hours "N" will be 200 So, the equation then becomes 200 = 100 * e^k*3

OpenStudy (marigirl):

okay i am good so far with (0, No) .. if i am not mistaken

OpenStudy (danjs):

Then it says, the original number was doubled in 3 hours. The original number is No, from when t=0. So double that is 2*No, when t=3

OpenStudy (danjs):

\[\large N=N _{0}*e^{k*t}\] For t=3, you have double original number 2*No So if you put t=3 into that equation, you get 2*No \[\huge 2*N _{0}=N _{0}*e^{k*3}\]

OpenStudy (danjs):

that you can solve for k, divide both sides by No first \[\large 2=e^{3k}\] Take natural log of both sides \[\large \ln(2)=\ln e^{3k}\]

OpenStudy (danjs):

recall the property for natural log base e of e gives \[\large \ln e^a =a\] so you have \[\large \ln(2)=3k\] \[\large k=\frac{ \ln(2) }{ 3 }\]

OpenStudy (marigirl):

oh yup! got it :D Thank you so much

OpenStudy (danjs):

welcome